40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 12 03:52:51 2011, kiwamu, Update, SUS, PRM OSEM adjustment 
    Reply  Sat Aug 13 01:54:18 2011, kiwamu, Update, SUS, PRM OSEM adjustment part II before.pngafter.png
Message ID: 5219     Entry time: Sat Aug 13 01:54:18 2011     In reply to: 5203
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: PRM OSEM adjustment part II 

Adjustment of the PRM OSEMs are done. The coils turned out to be healthy.

The malfunction was fixed. It was because the UL OSEM was too deeply inserted and barely touching the AR surface of the mirror.

 

(OSEM adjustment)

 + Excited POS at 6.5 Hz with an amplitude of 3000 cnts by the LOCKIN oscillator.

 + Looked at the signal of each sensor in frequency domain.

 + Maximized the excitation peak for each of the four face OSEMs by rotating them.

 + Minimized the excitation peak in the SIDE signal by rotating it.

 + Adjusted the OSEM translational position so that they are in the midpoint of the OSEM range.

 

(POS sensitivity check)

From the view point of the matrix inversion, one thing we want to have is the equally sensitive face sensors and insensitive SIDE OSEM to the POS motion.

To check the success level of today's PRM adjusment, I ran swept sine measurements to take the transfer function from POS to each sensor.

The plots below are the results.  The first figure is the one measured before the adjustment and the second plot is the one after the adjustment.

As shown in the plot, before the adjustment the sensitivity of OSEMs were very different and the SIDE OSEM is quite sensitive to the POS motion.

So PRM used be in an extremely bad situation.

After the adjustment, the plot became much better.

The four face sensors have almost the same sensitivity (within factor of 3) and the SIDE is quite insensitive to the POS motion.

before.png after.png

Quote from #5203

 We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.

The UL and LL coils are not working correctly, the forces are weak.

ELOG V3.1.3-